### v1

@@ Abstract.
It is well-known in practice, that $\mathbf{L}^{1}$ data fitting leads to improved robustness compared to standard $\mathbf{L}^{2}$ data fitting. However, it is unclear whether resulting algorithms will perform as well in case of regular data without outliers. In this paper, we therefore analyze generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity for Inverse Problems $F(u)=g$ in a general setting, including general measurement errors and errors in the forward operator. The derived results are then applied to the situation of discretized Gaussian white noise, and we show that the resulting error bounds allow for order-optimal rates of convergence. These findings are also investigated in numerical simulations.
@@ 1. Introduction
with an operator $F:D(F)\subset\mathcal{X}\to\mathbf{L}^{1}\left(\mathbb{M}\right)$ mapping between a Banach space $\mathcal{X}$ and $\mathbf{L}^{1}\left(\mathbb{M}\right)$, where $\mathbb{M}\subset\mathbb{R}^{d}$ is some smooth bounded domain. The exact solution of (1.1) will be denoted by $u^{\dagger}\in\mathcal{X}$, and the corresponding exact datum by $g^{\dagger}:=F(u^{\dagger})\in\mathbf{L}^{1}\left(\mathbb{M}\right)$. As typical in applications, we will assume that $g^{\dagger}$ is not (exactly) available, but rather some approximation $g^{\mathrm{obs}}\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$ of it, where $\left(\mathbb{M}_{h},\Sigma,\nu\right)$ is a (possibly discrete) finite measure space. Note, that this allows $g^{\mathrm{obs}}$ to live on a different domain than $g^{\dagger}$, e.g. for discretization reasons. This setup also allows for finite-dimensional data in the sense that $g^{\mathrm{obs}}$ can be interpreted as a vector in $\mathbb{R}^{n}$, but also for different types of infinite-dimensional data. For the moment we do not specify a model for $g^{\mathrm{obs}}$. Additionally, we will assume that $F$ is cannot be evaluated or is not known exactly, but only some approximation
@@
The rationale behind the choice of the $\mathbf{L}^{1}$ data fidelity term $\left\|\cdot\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)}$ in (1.2) is, that compared to the standard least-squares choice $\left\|\cdot\right\|_{\mathbf{L}^{2}\left(\mathbb{M}_{h}\right)}^{2}$, it is somewhat more robust w.r.t. outliers in the data. If e.g. $y_{1},\ldots,y_{n}\in\mathbb{R}$ are observations, then the minimizer $\bar{y}:=\operatorname*{argmin}_{y\in\mathbb{R}}\sum_{i=1}^{n}\left(y-y_{i}\right)^{2}$ is the empirical mean $\bar{y}=n^{-1}\sum_{i=1}^{n}y_{i}$. If one of the observations is far of the others, then this influences $\bar{y}$ with a weight $n^{-1}$. A more robust estimator is the median, which is given by $\tilde{y}:=\operatorname*{argmin}_{y\in\mathbb{R}}\sum_{i=1}^{n}\left|y-y_{i}\right|$, and is well-known to be independent of the specific value of the smallest or largest $y_{i}$, since it can practically be computed as the $1/2$-quantile of $y_{1},\ldots,y_{n}$. As a consequence, a better choice as data fidelity term in case of possible outliers in the data $g^{\mathrm{obs}}$ is $\left\|\cdot\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)}$ as in (1.2). This yields arguably a robust method, see e.g. [HW14, KWH16, CJ12] and the references therein.
@@
Generalized Tikhonov regularization, i.e. methods of the form (1.2), have been studied intensively over the last decades, see e.g. [Egg93, BO04, Res05, HKPS07, BKM+08, Gra10, BB11, WH12, Fle18a, KR19, BBKS20, HP20, PH25] to name a few. Investigated questions include existence of minimizers, possible uniqueness, stability w.r.t. $g^{\mathrm{obs}}$, and finally also convergence $\hat{u}_{\alpha}\to u^{\dagger}$ if $g^{\mathrm{obs}}\to g^{\dagger}$, $F_{h}\to F$ and $\alpha\searrow 0$ in a suitable sense.
@@
In this paper, we will mostly be interested in rates of convergence, this is quantitative estimates of $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$ for some loss function $\ell:\mathcal{X}\times\mathcal{X}\to\left[0,\infty\right)$ in terms of $\alpha$, the noise contained in $g^{\mathrm{obs}}$, the approximation parameter $h$ of $F_{h}$, and smoothness properties of $u^{\dagger}$. Since our final model for $g^{\mathrm{obs}}$ will be random, we are also interested in the expectation or concentration of $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$. Once such error bounds for $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$ are derived, a natural question is the optimality of those, i.e. whether methods different from (1.2) such as using $\mathbf{L}^{2}$ data fitting instead or different proof techniques could lead to improved estimates.
@@
For many specific classes of operators and noise models, the optimal rates of convergence are known, see e.g. [BHMR07, Cav08, RSH16, WSH20]. Especially, if the noise is white Gaussian with variance $\sigma^{2}>0$, and if the operator is a linear isomorphism between $H^{t}\left(\Omega\right)$ and $H^{t+a}\left(\mathbb{M}\right)$ for all $t\in\mathbb{R}$ with some fixed $a>0$, where $\Omega\subset\mathbb{R}^{m}$ is another smooth bounded domain, then under a source condition $u^{\dagger}\in H^{s}\left(\Omega\right)$ with $0<s<a$, the optimal rate of convergence is
@@
see e.g. [WSH20, Thm. 4.10].
In this paper, we aim to provide a new analysis for generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity (1.2). If $g^{\mathrm{obs}}$ is given as $g^{\dagger}$ plus impulsive noise (see the definition in Section 3 below), which is a continuous model for noise containing arbitrarily heavy outliers, this has been analyzed in [HW14, KWH16, HW22]. We build upon those results and aim to answer what happens if (1.2) is used in a discretized observation scheme. We will show that an extension of the previously mentioned analysis can be used to derive rates of convergence, which are known to be optimal in the Gaussian case. We also discuss efficient numerical methods for the minimization of (1.2) and investigate their performance in simulations. As a consequence, we recommend to use $\mathbf{L}^{1}$ data fitting as the standard method, since it has nearly no drawbacks compared to $\mathbf{L}^{2}$ data fitting, but provides a way more robust algorithm.
The remainder of this paper is organized as follows. In Section 2 we recall central results on generalized Tikhonov regularization (1.2), especially the tools required to derive error bounds as presented in [HW22]. The successional Section 3 is devoted to impulsive noise and provides the specific model introduced in [HW14, KWH16] and the resulting convergence rates in this case. In Section 4 we discuss the more realistic setting of discrete noise and derive error bounds for $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$, especially in case of discretized white Gaussian noise. The resulting rates of convergence turn out to equal (1.3), i.e. are known to be optimal in this case. Section 5 presents numerical simulations, which support the previous estimates. In Section 6 we briefly discuss our results and conclude.
@@ 2.1. Effective noise level
The rationale behind is, that we would like to use $u\mapsto\left\|F(u)-g^{\dagger}\right\|_{\mathbf{L}^{1}\left(\mathbb{M}\right)}$ as a data fidelity, i.e. to compute
@@
corresponding to the exact operator and noise free data, but this is not possible due to lack of knowledge. The definition of $\mathrm{err}$ implies that
@@ 2.2. Variational source condition
The error contribution of (ideal) generalized Tikhonov regularization (2.1) can be controlled as follows. Suppose we want to investigate the error w.r.t. some loss function $\ell:\mathcal{X}\times\mathcal{X}\to\left[0,\infty\right]$. Most common examples of losses are norm powers $\ell\left(u,u^{\dagger}\right)=\left\|u-u^{\dagger}\right\|_{\mathcal{X}}^{r}$ or Bregman divergences
@@
where $u^{*}\in\partial R\left(u^{\dagger}\right)$ denotes a subgradient. Now let us assume that $u^{\dagger}\in D(F)$ satisfies a variational source condition
@@
with $\varphi(t)=C\rho^{\frac{2a}{a+s}}t^{\frac{s}{a+s}},t>0$, for some constant $C>0$ independent of $\rho$ and $t$, see Example 2.2 in [HW22]. In this case, it follows via differentiation that
@@ 2.3. Bounding the effective noise level
for the effective noise level functional, corresponding to the fact that especially for random and / or operator noise, this functional will not be uniformly bounded for all $u\in D(F)$. In (2.5), $\mathcal{Y}_{\mathrm{d}}^{m}\subset\mathbf{L}^{1}\left(\mathbb{M}\right)$ for $m=1,\ldots,M$ are Banach spaces with $R(F)\subset\mathcal{Y}_{\mathrm{d}}^{m}$, and $\eta_{0},\ldots,\eta_{M}\geq 0$ are random variables (or constants in the deterministic case). Any specific model for $g^{\mathrm{obs}}$ and $F_{h}$ in (1.1) satisfying (2.5) is called $\left(\mathcal{Y}_{\mathrm{d}}^{1},\ldots,\mathcal{Y}_{\mathrm{d}}^{M}\right)$-admissible.
@@
for all $\delta\in[0,\delta_{0}]$, where $\gamma:[0,\delta_{0}]\to[0,\infty)$ is a function and $\delta_{0}>0$, $r\geq 1$ are constants. The parameter $r$ is only to take care of potential exponents in $\ell$ e.g. in case of norm losses. The bound (2.6), if assumed for each space $\mathcal{Y}_{\mathrm{d}}^{m}$, $m=1,\ldots,M$, allows to further estimate each of the additional terms in (2.5) separately.
@@
Suppose $\mathbb{M}\subset\mathbb{R}^{d}$ is a bounded Lipschitz domain and $F:D(F)\subset X\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ for $k>\frac{d}{p}$ with the standard $\mathbf{L}^{p}$-based Sobolev space $\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ is Lipschitz continuous in the sense that
(2.7) $\left\|F(u)-F\left(u^{\dagger}\right)\right\|_{\mathbf{W}^{k,p}\left(\mathbb{M}\right)}\leq L\ell\left(u,u^{\dagger}\right)^{\frac{1}{r}}.$
Then $F$ is $\left(\mathbf{L}^{\infty}\left(\mathbb{M}\right),\gamma,\ell\right)$-smoothing with $\gamma\left(\delta\right)=C\delta^{\frac{k}{d}-\frac{1}{p}}$ for some constant $C>0$ independent of $\delta$.
@@
where $a>0$ is again a parameter, then - for $a\in\mathbb{N}$ and $\ell=\left\|\cdot-\cdot\right\|_{\mathbf{L}^{2}\left(\Omega\right)}$ - this implies (2.7) with $k=a$ and $p=2$. In this sense, (2.7) can be seen as an at least smoothing assumption compared to the at most smoothing assumption (2.3).
@@ 2.4. Error bounds
Suppose the data model for $g^{\mathrm{obs}}$ is $\left(\mathcal{Y}_{\mathrm{d}}^{1},\ldots,\mathcal{Y}_{\mathrm{d}}^{M}\right)$-admissible with $\eta_{0},\ldots,\eta_{M}\geq 0$ as in (2.5) and that the variational source condition (2.2) is satisfied. For those $m\in\{1,\ldots,M\}$ with $\eta_{m}>0$, we additionally assume that $F$ is $\left(\mathcal{Y}_{\mathrm{d}}^{m},\gamma_{m},\ell\right)$-smoothing as in (2.6). If $4\eta_{m}<\delta_{0}^{m}$ for all such $m$, then each global minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) satisfies

### v2

@@ Abstract.
It is well-known in practice that $\mathbf{L}^{1}$ data fitting leads to improved robustness compared to standard $\mathbf{L}^{2}$ data fitting. However, it is unclear whether resulting algorithms will perform as well in case of regular data without outliers. In this paper, we therefore analyze generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity for Inverse Problems $F(u)=g$ in a general setting, including general measurement errors and errors in the forward operator. The derived results are then applied to the situation of discretized Gaussian white noise, and we show that the resulting error bounds allow for order-optimal rates of convergence. These findings are also investigated in numerical simulations.
@@ 1. Introduction
with an operator $F:D(F)\subset\mathcal{X}\to\mathbf{L}^{1}\left(\mathbb{M}\right)$ mapping between a Banach space $\mathcal{X}$ and $\mathbf{L}^{1}\left(\mathbb{M}\right)$, where $\mathbb{M}\subset\mathbb{R}^{d}$ is some smooth bounded domain. The exact solution of (1.1) will be denoted by $u^{\dagger}\in\mathcal{X}$, and the corresponding exact datum by $g^{\dagger}:=F(u^{\dagger})\in\mathbf{L}^{1}\left(\mathbb{M}\right)$. As typical in applications, we will assume that $g^{\dagger}$ is not (exactly) available, but rather some approximation $g^{\mathrm{obs}}\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$ of it, where $\left(\mathbb{M}_{h},\Sigma,\nu\right)$ is a (possibly discrete) finite measure space. Note that this allows $g^{\mathrm{obs}}$ to live on a different domain than $g^{\dagger}$, e.g. for discretization reasons. This setup also allows for finite-dimensional data in the sense that $g^{\mathrm{obs}}$ can be interpreted as a vector in $\mathbb{R}^{n}$, but also for different types of infinite-dimensional data. For the moment we do not specify a model for $g^{\mathrm{obs}}$. Additionally, we will assume that $F$ cannot be evaluated or is not known exactly, but only some approximation
@@
The rationale behind the choice of the $\mathbf{L}^{1}$ data fidelity term $\left\|\cdot\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)}$ in (1.2) is, that compared to the standard least-squares choice $\left\|\cdot\right\|_{\mathbf{L}^{2}\left(\mathbb{M}_{h}\right)}^{2}$, it is somewhat more robust w.r.t. outliers in the data. If e.g. $y_{1},\ldots,y_{n}\in\mathbb{R}$ are observations, then the minimizer $\bar{y}:=\operatorname*{argmin}_{y\in\mathbb{R}}\sum_{i=1}^{n}\left(y-y_{i}\right)^{2}$ is the empirical mean $\bar{y}=n^{-1}\sum_{i=1}^{n}y_{i}$. If one of the observations is far off the others, then this influences $\bar{y}$ with a weight $n^{-1}$. A more robust estimator is the median, which is given by $\tilde{y}:=\operatorname*{argmin}_{y\in\mathbb{R}}\sum_{i=1}^{n}\left|y-y_{i}\right|$, and is well-known to be independent of the specific value of the smallest or largest $y_{i}$, since it can practically be computed as the $1/2$-quantile of $y_{1},\ldots,y_{n}$. As a consequence, a better choice as data fidelity term in case of possible outliers in the data $g^{\mathrm{obs}}$ is $\left\|\cdot\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)}$ as in (1.2). This yields arguably a robust method, see e.g. [HW14, KWH16, CJ12] and the references therein.
@@
Generalized Tikhonov regularization, i.e. methods of the form (1.2), has been studied intensively over the last decades, see e.g. [Egg93, BO04, Res05, HKPS07, BKM+08, Gra10, BB11, WH12, Fle18a, KR19, BBKS20, HP20, PH25] to name a few. Investigated questions include existence of minimizers, possible uniqueness, stability w.r.t. $g^{\mathrm{obs}}$, and finally also convergence $\hat{u}_{\alpha}\to u^{\dagger}$ if $g^{\mathrm{obs}}\to g^{\dagger}$, $F_{h}\to F$ and $\alpha\searrow 0$ in a suitable sense.
@@
In this paper, we will mostly be interested in rates of convergence, that is, quantitative estimates of $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$ for some loss function $\ell:\mathcal{X}\times\mathcal{X}\to\left[0,\infty\right]$ in terms of $\alpha$, the noise contained in $g^{\mathrm{obs}}$, the approximation parameter $h$ of $F_{h}$, and smoothness properties of $u^{\dagger}$. Since our final model for $g^{\mathrm{obs}}$ will be random, we are also interested in the expectation or concentration of $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$. Once such error bounds for $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$ are derived, a natural question is the optimality of those, i.e. whether methods different from (1.2) such as using $\mathbf{L}^{2}$ data fitting instead or different proof techniques could lead to improved estimates.
@@
For many specific classes of operators $F:D(F)\subseteq\mathbf{L}^{2}\left(\Omega\right)\to\mathbf{L}^{2}\left(\mathbb{M}\right)$ and noise models, the optimal rates of convergence are known, see e.g. [BHMR07, Cav08, RSH16, WSH20]. Especially, if the noise is white Gaussian with variance $\sigma^{2}>0$, and if the operator is a linear isomorphism between $H^{t}\left(\Omega\right)$ and $H^{t+a}\left(\mathbb{M}\right)$ for all $t\in\mathbb{R}$ with some fixed $a>0$, where $\Omega\subset\mathbb{R}^{m}$ is another smooth bounded domain, then under a source condition $u^{\dagger}\in H^{s}\left(\Omega\right)$ with $0<s<a$, the optimal rate of convergence is
@@
see e.g. [WSH20, Thm. 4.10]. Note that any such operator can, by boundedness of $\mathbb{M}$ and hence $\mathbf{L}^{\infty}\left(\mathbb{M}\right)\hookrightarrow\mathbf{L}^{1}\left(\mathbb{M}\right)$, also be considered as an operator mapping into $\mathbf{L}^{1}\left(\mathbb{M}\right)$ as soon as $a>d/2$ for $\mathcal{X}=\mathbf{L}^{2}\left(\Omega\right)$. This is the case which will be considered mostly below.
In this paper, we aim to provide a new analysis for generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity (1.2). If $g^{\mathrm{obs}}$ is given as $g^{\dagger}$ plus impulsive noise (see the definition in Section 3 below), which is a continuous model for noise containing arbitrarily heavy outliers, this has been analyzed in [HW14, KWH16, HW22]. We build upon those results and aim to answer what happens if (1.2) is used in a discretized observation scheme. We will show that an extension of the previously mentioned analysis can be used to derive rates of convergence, which are known to be optimal in the Gaussian case. We also discuss efficient numerical methods for the minimization of (1.2) and investigate their performance in simulations. As a consequence, we recommend to use $\mathbf{L}^{1}$ data fitting as the standard method, since it has nearly no drawbacks compared to $\mathbf{L}^{2}$ data fitting, but provides a significantly more robust algorithm.
The remainder of this paper is organized as follows. In Section 2 we recall central results on generalized Tikhonov regularization (1.2), especially the tools required to derive error bounds as presented in [HW22]. The successive Section 3 is devoted to impulsive noise and provides the specific model introduced in [HW14, KWH16] and the resulting convergence rates in this case. In Section 4 we discuss the more realistic setting of discrete noise and derive error bounds for $\ell\left(\hat{u}_{\alpha},u^{\dagger}\right)$, especially in case of discretized white Gaussian noise. The resulting rates of convergence turn out to equal (1.3), i.e. are known to be optimal in this case. Section 5 presents numerical simulations, which support the previous estimates. In Section 6 we briefly discuss our results and conclude.
@@ 2.1. Effective noise level
The rationale behind is that we would like to use $u\mapsto\left\|F(u)-g^{\dagger}\right\|_{\mathbf{L}^{1}\left(\mathbb{M}\right)}$ as a data fidelity, i.e. to compute
@@
corresponding to the exact operator and noise-free data, but this is not possible due to lack of knowledge. The definition of $\mathrm{err}$ implies that
@@ 2.2. Variational source condition
The error contribution of (ideal) generalized Tikhonov regularization (2.1) can be controlled as follows. Suppose we want to investigate the error w.r.t. some loss function $\ell:\mathcal{X}\times\mathcal{X}\to\left[0,\infty\right]$. The most common examples of losses are norm powers $\ell\left(u,u^{\dagger}\right)=\left\|u-u^{\dagger}\right\|_{\mathcal{X}}^{r}$ or Bregman divergences
@@
where $u^{*}\in\partial\mathcal{R}\left(u^{\dagger}\right)$ denotes an element of the subdifferential $\partial\mathcal{R}\left(u^{\dagger}\right)$. Now let us assume that $u^{\dagger}\in D(F)$ satisfies a variational source condition
@@
with $\varphi(t)=C\rho^{\frac{2a}{a+s}}t^{\frac{s}{a+s}},t>0$, for some constant $C>0$ independent of $\rho$ and $t$, see Example 2.2 in [HW22]. In this case, it follows via convex conjugation properties that
@@ 2.3. Bounding the effective noise level
for the effective noise level functional, reflecting the fact that, especially for random and / or operator noise, this functional will not be uniformly bounded for all $u\in D(F)$. In (2.5), $\mathcal{Y}_{\mathrm{d}}^{m}\subset\mathbf{L}^{1}\left(\mathbb{M}\right)$ for $m=1,\ldots,M$ are Banach spaces with $R(F)\subset\mathcal{Y}_{\mathrm{d}}^{m}$, and $\eta_{0},\ldots,\eta_{M}\geq 0$ are random variables (or constants in the deterministic case). Any specific model for $g^{\mathrm{obs}}$ and $F_{h}$ in (1.1) satisfying (2.5) is called $\left(\mathcal{Y}_{\mathrm{d}}^{1},\ldots,\mathcal{Y}_{\mathrm{d}}^{M}\right)$-admissible.
@@
for all $\delta\in[0,\delta_{0}]$, where $\gamma:[0,\delta_{0}]\to[0,\infty)$ is a function and $\delta_{0}>0$, $r\geq 1$ are constants. The parameter $r$ is only to take care of potential exponents in $\ell$ e.g. in case of norm losses. The bound (2.6), if assumed for each space $\mathcal{Y}_{\mathrm{d}}^{m}$, $m=1,\ldots,M$, allows us to further estimate each of the additional terms in (2.5) separately.
@@
Suppose $\mathbb{M}\subset\mathbb{R}^{d}$ is a bounded Lipschitz domain and $F:D(F)\subset X\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ is Lipschitz continuous in the sense that
(2.7) $\left\|F(u)-F\left(u^{\dagger}\right)\right\|_{\mathbf{W}^{k,p}\left(\mathbb{M}\right)}\leq L\ell\left(u,u^{\dagger}\right)^{\frac{1}{r}},$
where $k>\frac{d}{p}$ and $\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ denotes the standard $\mathbf{L}^{p}$-based Sobolev space. Then $F$ is $\left(\mathbf{L}^{\infty}\left(\mathbb{M}\right),\gamma,\ell\right)$-smoothing with $\gamma\left(\delta\right)=C\delta^{\frac{k}{d}-\frac{1}{p}}$ for some constant $C>0$ independent of $\delta$.
@@
where $a>0$ is again a parameter, then, for $a\in\mathbb{N}$ and $\ell=\left\|\cdot-\cdot\right\|_{\mathbf{L}^{2}\left(\Omega\right)}$, this implies (2.7) with $k=a$ and $p=2$. In this sense, (2.7) can be seen as an at least smoothing assumption compared to the at most smoothing assumption (2.3).
@@ 2.4. Error bounds
Suppose the data model for $g^{\mathrm{obs}}$ is $\left(\mathcal{Y}_{\mathrm{d}}^{1},\ldots,\mathcal{Y}_{\mathrm{d}}^{M}\right)$-admissible with $\eta_{0},\ldots,\eta_{M}\geq 0$ as in (2.5) and that the variational source condition (2.2) is satisfied. Suppose furthermore that $r>1$. For those $m\in\{1,\ldots,M\}$ with $\eta_{m}>0$, we additionally assume that $F$ is $\left(\mathcal{Y}_{\mathrm{d}}^{m},\gamma_{m},\ell\right)$-smoothing as in (2.6). If $4\eta_{m}<\delta_{0}^{m}$ for all such $m$, then each global minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) satisfies